# The Complexity of Resilience Problems via Valued Constraint Satisfaction

### v5

@@ 8.4. Infinite Duals
Note that items (8)-(10) are just cyclic shifts of the same condition. Again one can verify that ${\mathfrak{N}}\models\neg\mu$, because ${\mathfrak{C}}_{\mu}\models\neg\mu$. Therefore, there is an embedding $g$ of ${\mathfrak{N}}$ into ${\mathfrak{C}}_{\mu}$. By the definition of ${\mathfrak{N}}$, the operation $g$ is pseudo cyclic and canonical with respect to $\operatorname{Aut}({\mathfrak{C}}_{\mu})$; it has the behavior of a minority operation on orbits of pairs of distinct elements.
@@
We remark that there is no binary fractional polymorphism of $\Gamma$ that is canonical and pseudo cyclic with respect to $\operatorname{Aut}({\mathfrak{C}}_{\mu})$; this is shown in detail in (Žaneta Semanišinová, 2025, Example 5.23).
@@ 8.5. The Resilience Tractability Conjecture
Let $\mu$ be a union of connected conjunctive queries over the signature $\tau$, and let $\sigma\subseteq\tau$. If the structure ($(\{0,1\};\operatorname{OIT})$ has no pp-construction in $\Gamma_{\mu,\sigma}$, then there exists a dual ${\mathfrak{B}}$ of $\mu$ such that $\operatorname{Aut}({\mathfrak{B}})$ contains an automorphism group of a finitely bounded homogeneous structure ${\mathfrak{A}}$ and $\Gamma:=\Gamma({\mathfrak{B}},\sigma)$ has a fractional polymorphism of arity $\ell\geq 2$ which is canonical and pseudo cyclic with respect to $\operatorname{Aut}({\mathfrak{A}})$ (and in this case, $\operatorname{VCSP}(\Gamma)$ is in P by Theorem 7.18).
@@
For the following conjunctive query $\mu$, the NP-hardness of the resilience problem without multiplicities was shown in (Freire et al., 2015b); to illustrate our condition, we verify that the structure ($(\{0,1\};\operatorname{OIT})$ has a pp-construction in $\Gamma_{\mu}$ and thus prove in a different way that the resilience problem (with multiplicities) for $\mu$ is NP-hard.
@@
The resilience problem without multiplicities for $\mu$ is NP-complete (Freire et al., 2015b), and hence $\operatorname{VCSP}(\Gamma_{\mu})$ is NP-hard (Proposition 8.20). Since the Gaifman graph of $\mu$ is NP-complete, the structure ${\mathfrak{C}}_{\mu}$ from Theorem 8.19 exists. Let $\Gamma:=\Gamma({\mathfrak{C}}_{\mu},\emptyset)$. We provide a pp-construction of ($(\{0,1\};\operatorname{OIT})$ in $\Gamma$, which also proves NP-hardness of $\operatorname{VCSP}(\Gamma)$ and hence the resilience problem of $\mu$ with multiplicities by Corollary 5.14. Since $\Gamma$ is homomorphically equivalent to $\Gamma_{\mu}$, this also provides a pp-construction of ($(\{0,1\};\operatorname{OIT})$ in $\Gamma_{\mu}$ (see Lemma 5.16).
Let $C$ be the domain of $\Gamma$. Denote the relations $\operatorname{Opt}(R)$, $\operatorname{Opt}(S)$, $\operatorname{Opt}(T)$ by $R^{*}$, $S^{*}$, $T^{*}$, respectively. In the following, for $U\in\{R,S,T\}$ and variables $x,y$ we write $2U(x,y)$ for short instead of $U(x,y)+U(x,y)$. Let $\phi(a,b,c,d,e,f,g,h,i)$ be the expression
@@
Figure 3. Example 8.25, visualisation of $\mu$ and $\phi$. The thick edges cannot be removed.
[…]
Arrow depiction of the query and the expression from the example, with thick arrows symbolizing the crisp constraints.
In what follows, we say that an atomic $\tau$-expression holds if it evaluates to 00 and an atomic $\tau$-expression is violated if it does not hold. Since there are $7$ non-overlapping copies of $\mu$ in $\phi$, the cost of $\phi$ is at least $7$. Every assignment where
@@
where the reverse implication uses that ${\mathfrak{C}}_{\mu}$ is homogeneous and embeds all finite structures that do not satisfy $\mu$. Define $RS\in\langle\Gamma\rangle$ by
@@
Then $\Delta$ is homomorphically equivalent to ($(\{0,1\};\operatorname{OIT})$, witnessed by the homomorphism from $\Delta$ to ($(\{0,1\};\operatorname{OIT})$ that maps $(u,v)$ to $1$ if $R(u,v)$ and to 00 otherwise, and the homomorphism ($(\{0,1\};\operatorname{OIT})\to\Delta$ that maps $1$ to any pair of vertices $(u,v)\in R$ and 00 to any pair of vertices $(u,v)\notin R$. Therefore, $\Gamma$ pp-constructs ($(\{0,1\};\operatorname{OIT})$.
@@ 8.6. An example of formerly open complexity
Let $\tau=\{R,S\}$ be the signature of $\mu$. To study the complexity of resilience of $\mu$, it will be convenient to work with a dual which has different model-theoretic properties than the duals ${\mathfrak{B}}_{\mu}$ from Theorem 8.18 and ${\mathfrak{C}}_{\mu}$ from Theorem 8.19, namely a dual that is a model-complete core.
@@
Note that the definition of cores of valued structures with finite domain (Definition 7.12) and the definition above specialise to the same concept for relational structures over finite domains. A structure with an oligomorphic automorphism group is a model-complete core if and only if for every $n\in{\mathbb{N}}$ every orbit of $n$-tuples can be defined with an existential positive formula (Bodirsky, 2021). Every countable structure ${\mathfrak{B}}$ is homomorphically equivalent to a model-complete core, which is unique up to isomorphism (Bodirsky, 2007, 2021); we refer to this structure as the model-complete core of ${\mathfrak{B}}$. The advantage of working with model-complete cores is that the structure is in a sense ‘minimal’ and therefore easier to work with in concrete examples. [^3] The model-complete core of ${\mathfrak{B}}_{\mu}$ would be a natural choice for the canonical dual of $\mu$ to work with instead of ${\mathfrak{B}}_{\mu}$. However, proving that the model-complete core has a finitely bounded homogeneous expansion (so that, for example, Theorem 3.4 applies) requires introducing further model-theoretical notions (Mottet and Pinsker, 2021) which we want to avoid in this article.
Figure 4. Visualisation of the query $\mu$ from (18).
[…]
Arrow depiction of the query.
@@
There is a finitely bounded homogeneous dual ${\mathfrak{B}}$ of $\mu$ such that the valued $\tau$-structure $\Gamma:=\Gamma({\mathfrak{B}},\emptyset)$ has a binary fractional polymorphism which is canonical and pseudo cyclic with respect to $\operatorname{Aut}({\mathfrak{B}})=\operatorname{Aut}(\Gamma)$. Hence, $\operatorname{VCSP}(\Gamma)$ and the resilience problem for $\mu$ are in P. The polynomial-time tractability result even holds for resilience of $\mu$ with exogeneous relations from any $\sigma\subseteq\tau$.
Proof.
Figure 5. Illustration of a finite substructure of ${\mathfrak{B}}$ from the proof of Proposition 8.27 that contains representatives for all orbits of pairs of $\operatorname{Aut}({\mathfrak{B}})$. Arrows are not drawn on undirected edges.
[…]
Arrow depiction of all possible orbits of pairs of the automorphism group. Contains all possibilities of how relations S and R may be imposed on a pair of vertices.
Since the Gaifman graph of $\mu$ is a complete graph, there exists the structure ${\mathfrak{C}}_{\mu}$ as in Theorem 8.19. Let ${\mathfrak{B}}$ be the model-complete core of ${\mathfrak{C}}_{\mu}$. Note that ${\mathfrak{B}}$ has the property that a countable structure ${\mathfrak{A}}$ maps homomorphically to ${\mathfrak{B}}$ if and only if ${\mathfrak{A}}\models\neg\mu$; in particular, ${\mathfrak{B}}$ is a dual of $\mu$ and ${\mathfrak{B}}\models\neg\mu$. The structure ${\mathfrak{C}}_{\mu}$ is homogeneous, and it is known that the model-complete core of a homogeneous structure is again homogeneous (see Proposition 4.7.7 in (Bodirsky, 2021)), so ${\mathfrak{B}}$ is homogeneous. Let $\Gamma:=\Gamma({\mathfrak{B}},\emptyset)$.
@@
Suppose for contradiction that (21) does not hold for some distinct $x$ and $y$. Then $\neg S(x)\vee\neg R(y,y)$ and $\neg R(x,x)\vee\neg S(y)$, i.e., $\neg S(x)\wedge\neg R(x,x)$, or $\neg S(x)\wedge\neg S(y)$, or $\neg R(y,y)\wedge\neg R(x,x)$, or $\neg R(y,y)\wedge\neg S(y)$. In each of these cases we may add both $R$-edges between the distinct elements $x$ and $y$ to ${\mathfrak{B}}$ and obtain a structure not satisfying $\mu$, which leads to a contradiction as above.
For an illustration of a finite substructure of ${\mathfrak{B}}$ which contains a representative for every orbit of pairs in $\operatorname{Aut}({\mathfrak{B}})$, see Figure 5.
Claim 1.
For every finite $\tau$-structure ${\mathfrak{A}}$ that satisfies $\neg\mu$ and the sentences in (20) and (21), there exists a strong homomorphism to ${\mathfrak{B}}$, i.e., a homomorphism that also preserves the complements of $R$ and $S$. First observe that ${\mathfrak{B}}$ embeds the relational $\tau$-structure ${\mathfrak{A}}^{\prime}$ on the domain $A^{\prime}=\mathbb{N}\cup\{a\}\cup\{b\}$, where $a,b\notin\mathbb{N}$ are distinct, $S^{{\mathfrak{A}}^{\prime}}=\{a\}$ and
@@
This is because ${\mathfrak{A}}^{\prime}\models\neg\mu$ and hence ${\mathfrak{A}}^{\prime}$ maps homomorphically to ${\mathfrak{B}}$ and unless embedded, it contradicts ${\mathfrak{B}}\not\models\mu$. In particular, there are infinitely many $x\in B$ such that ${\mathfrak{B}}\models\neg S(x)\wedge\neg R(x,x)$ and by (21), for every $y\in B$, $x\neq y$, we have ${\mathfrak{B}}\models R(x,y)\wedge R(y,x)$.
@@
Let $h$ be a homomorphism from ${\mathfrak{A}}$ to ${\mathfrak{B}}$, which exists since ${\mathfrak{A}}\models\neg\mu$. If $s(h)+r(h)=0$, then $h$ is a strong homomorphism and there is nothing to prove. Suppose therefore $s(h)+r(h)>0$. We construct a homomorphism $h^{\prime}$ such that $r(h^{\prime})+s(h^{\prime})<r(h)+s(h)$. Since $r(h)+s(h)$ is finite, by applying this construction finitely many times, we obtain a strong homomorphism from ${\mathfrak{A}}$ to ${\mathfrak{B}}$.
If $s(h)>0$, then there exists $a\in A\setminus S^{{\mathfrak{A}}}$ such that $h(a)\in S^{{\mathfrak{B}}}$. By (19), ${\mathfrak{B}}\not\models R(h(a),h(a))$ and hence ${\mathfrak{A}}\not\models R(a,a)$. Pick $b\in B\setminus h(A)$ such that ${\mathfrak{B}}\models\neg S(b)\wedge\neg R(b,b)$ and define
@@
Observe that $h^{\prime}$ is a homomorphism, $s(h^{\prime})<s(h)$ and $r(h^{\prime})=r(h)$. If $r(h)>0$, then there exists $(x,y)\in A^{2}\setminus R^{{\mathfrak{A}}}$ such that $(h(x),h(y))\in R^{{\mathfrak{B}}}$. If $x=y$, the argument is similar as in the case $s(h)>0$. Finally, if $x\neq y$, then ${\mathfrak{A}}\models(S(x)\wedge R(y,y))\vee(R(x,x)\wedge S(y))$, because ${\mathfrak{A}}$ satisfies the sentence in (21). Since ${\mathfrak{A}}$ satisfies the sentence in (20), ${\mathfrak{A}}\models R(y,x)$. Since $h$ is a homomorphism, we have
@@
which contradicts ${\mathfrak{B}}\not\models\mu$.
Claim 2.
Every finite $\tau$-structure ${\mathfrak{A}}$ that satisfies $\neg\mu$ and the universal sentences in (20) and (21) embeds into ${\mathfrak{B}}$. In particular, ${\mathfrak{B}}$ is finitely bounded. Let ${\mathfrak{A}}$ be such a structure. By Theorem 8.19, there is an embedding $e$ of ${\mathfrak{A}}$ into ${\mathfrak{C}}_{\mu}$. Since ${\mathfrak{C}}_{\mu}$ is homogeneous and embeds every finite $\tau$-structure that satisfies $\neg\mu$, there exists a finite substructure ${\mathfrak{A}}^{\prime}$ of ${\mathfrak{C}}_{\mu}$ satisfying the sentences in (20) and (21) such that $e({\mathfrak{A}})$ is a substructure of ${\mathfrak{A}}^{\prime}$ and for all distinct $a,b\in A$ there exists $s\in S^{{\mathfrak{A}}^{\prime}}$ such that ${\mathfrak{C}}_{\mu}\models R(e(a),s)\wedge R(s,e(b))$. By Claim 1, there is a strong homomorphism $h$ from ${\mathfrak{A}}^{\prime}$ to ${\mathfrak{B}}$.
We claim that $h\circ e$ is injective and therefore an embedding of ${\mathfrak{A}}$ into ${\mathfrak{B}}$. Suppose there exist distinct $a,b\in A$ such that $h(e(a))=h(e(b))$. Since $e({\mathfrak{A}})$ satisfies the sentence in (20), we have ${\mathfrak{B}}\models R(h(e(a)),h(e(a)))$. Let $s\in S^{{\mathfrak{A}}^{\prime}}$ be such that ${\mathfrak{C}}_{\mu}\models R(e(a),s)\wedge R(s,e(b))$. Hence,
@@
a contradiction to ${\mathfrak{B}}\not\models\mu$. It follows that $h\circ e$ is an embedding of ${\mathfrak{A}}$ into ${\mathfrak{B}}$.
@@
[…]
@@
Figure 6. An illustration of the conditions (A) and (B) in ${\mathfrak{M}}$ and ${\mathfrak{N}}$.
[…]
Arrow depiction of conditions (A)-(D), always showing the conditions for (x1, x2) on the horizontal axis, the conditions for (y1, y2) on the vertical axis and the resulting situation in M or N.
Claim 3.
${\mathfrak{M}}$ and ${\mathfrak{N}}$ satisfy the sentence in (20). We prove the statement for ${\mathfrak{M}}$; the proof for ${\mathfrak{N}}$ is similar. Let $(x_{1},y_{1}),(x_{2},y_{2})\in B$ be such that $(x_{1},y_{1})\neq(x_{2},y_{2})$ and ${\mathfrak{M}}\models\neg R((x_{2},y_{2}),(x_{1},y_{1}))$. Since ${\mathfrak{M}}$ satisfies the sentence in (21), we must have either that ${\mathfrak{M}}\models S(x_{1},y_{1})\wedge R((x_{2},y_{2}),(x_{2},y_{2}))$ or that ${\mathfrak{M}}\models S(x_{2},y_{2})\wedge R((x_{1},y_{1}),(x_{1},y_{1}))$. Suppose the former is true; the other case is treated analogously. Then ${\mathfrak{B}}\models R(x_{2},x_{2})\wedge R(y_{2},y_{2})$ and ${\mathfrak{B}}\models S(x_{1})\vee S(y_{1})$. If ${\mathfrak{B}}\models S(x_{1})$, then $x_{1}\neq x_{2}$ and by (20) we have ${\mathfrak{B}}\models R(x_{1},x_{2})\vee R(x_{2},x_{1})$. By (20) and (21) for $(y_{1},y_{2})$, we obtain that ${\mathfrak{M}}\models R((x_{1},y_{1}),(x_{2},y_{2}))$ by (22) or one of the conditions (A)-(D). The argument if ${\mathfrak{B}}\models S(y_{1})$ is similar with $x$ and $y$ switched.
Claim 4.
${\mathfrak{M}}$ and ${\mathfrak{N}}$ satisfy $\neg\mu$. Let $x_{1},x_{2},y_{1},y_{2}\in B$. Suppose for contradiction that

### v6

@@ 8.4. Infinite Duals
Note that items (8)-(10) are just cyclic shifts of the same condition. Again one can verify that ${\mathfrak{N}}\models\neg\mu$, because ${\mathfrak{H}}_{\mu}\models\neg\mu$. Therefore, there is an embedding $g$ of $\mathfrak{N}$} into ${\mathfrak{H}}_{\mu}$. By the definition of $\mathfrak{N}$}, the operation $g$ is pseudo cyclic and canonical with respect to $\operatorname{Aut}({\mathfrak{H}}_{\mu})$; it has the behavior of a minority operation on orbits of pairs of distinct elements.
@@
We remark that there is no binary fractional polymorphism of $\Gamma$ that is canonical and pseudo cyclic with respect to $\operatorname{Aut}({\mathfrak{H}}_{\mu})$; this is shown in detail in (Žaneta Semanišinová, 2025, Example 5.23).
@@ 8.5. The Resilience Tractability Conjecture
Let $\mu$ be a union of connected conjunctive queries over the signature $\tau$, and let $\sigma\subseteq\tau$. If the structure $(\{0,1\};\operatorname{OIT})$ has no pp-construction in $\Gamma_{\mu,\sigma}$, then there exists a dual $\mathfrak{B}$} of $\mu$ such that $\operatorname{Aut}({\mathfrak{B}})$ contains an automorphism group of a finitely bounded homogeneous structure $\mathfrak{A}$} and $\Gamma:=\Gamma({\mathfrak{B}},\sigma)$ has a fractional polymorphism of arity $\ell\geq 2$ which is canonical and pseudo cyclic with respect to $\operatorname{Aut}({\mathfrak{A}})$ (and in this case, $\operatorname{VCSP}(\Gamma)$ is in P by Theorem 7.19).
@@
For the following conjunctive query $\mu$, the NP-hardness of the resilience problem without multiplicities was shown in (Freire et al., 2015b); to illustrate our condition, we verify that the structure $(\{0,1\};\operatorname{OIT})$ has a pp-construction in $\Gamma_{\mu}$ and thus prove in a different way that the resilience problem (with multiplicities) for $\mu$ is NP-hard.
@@
The resilience problem without multiplicities for $\mu$ is NP-complete (Freire et al., 2015b), and hence $\operatorname{VCSP}(\Gamma_{\mu})$ is NP-hard (Proposition 8.20). Since the Gaifman graph of $\mu$ is NP-complete, the structure ${\mathfrak{H}}_{\mu}$ from Theorem 8.19 exists. Let $\Gamma:=\Gamma({\mathfrak{H}}_{\mu},\emptyset)$. We provide a pp-construction of $(\{0,1\};\operatorname{OIT})$ in $\Gamma$, which also proves NP-hardness of $\operatorname{VCSP}(\Gamma)$ and hence the resilience problem of $\mu$ with multiplicities by Corollary 5.17. Since $\Gamma$ is homomorphically equivalent to $\Gamma_{\mu}$, this also provides a pp-construction of $(\{0,1\};\operatorname{OIT})$ in $\Gamma_{\mu}$ (see Lemma 5.19).
Let $C$ be the domain of $\Gamma$. Denote the relations $\operatorname{Opt}(R)$, $\operatorname{Opt}(S)$, $\operatorname{Opt}(T)$ by $R^{*}$, $S^{*}$, $T^{*}$, respectively. In the following, for $U\in\{R,S,T\}$ and variables $x,y$ we write $2U(x,y)$ for short instead of $U(x,y)+U(x,y)$. Let $\phi(a,b,c,d,e,f,g,h,i)$ be the pp-expression
@@
Figure 4. Example 8.27, visualization of $\mu$ and $\phi$. The thick edges cannot be removed.:
In what follows, we say that an atomic $\tau$-expression holds if it evaluates to 0 and an atomic $\tau$-expression is violated if it does not hold. Since there are $7$ non-overlapping copies of $\mu$ in $\phi$, the cost of $\phi$ is at least $7$. Every assignment where
@@
where the reverse implication uses that ${\mathfrak{H}}_{\mu}$ is homogeneous and embeds all finite structures that do not satisfy $\mu$. Define $RS\in\langle\Gamma\rangle$ by
@@
Then $\Delta$ is homomorphically equivalent to $(\{0,1\};\operatorname{OIT})$, witnessed by the homomorphism from $\Delta$ to $(\{0,1\};\operatorname{OIT})$ that maps $(u,v)$ to $1$ if $R(u,v)$ and to 0 otherwise, and the homomorphism $(\{0,1\};\operatorname{OIT})\to\Delta$ that maps $1$ to any pair of vertices $(u,v)\in R$ and 0 to any pair of vertices $(u,v)\notin R$. Therefore, $\Gamma$ $p$-constructs $(\{0,1\};\operatorname{OIT})$.
@@ 8.6. An example of formerly open complexity
Let $\tau=\{R,S\}$ be the signature of $\mu$. To study the complexity of resilience of $\mu$, it will be convenient to work with a dual which has different model-theoretic properties than the duals ${\mathfrak{B}}_{\mu}$ from Theorem 8.18 and ${\mathfrak{H}}_{\mu}$ from Theorem 8.19, namely a dual that is a model-complete core.
@@
Note that the definition of cores of valued structures with finite domain (Definition 7.12) and the definition above specialize to the same concept for relational structures over finite domains. A structure with an oligomorphic automorphism group is a model-complete core if and only if for every $n\in{\mathbb{N}}$ every orbit of $n$-tuples can be defined with an existential positive formula (Bodirsky, 2021). Every countable structure $\mathfrak{B}$} is homomorphically equivalent to a model-complete core, which is unique up to isomorphism (Bodirsky, 2007, 2021); we refer to this structure as the model-complete core of $\mathfrak{B}$}. The advantage of working with model-complete cores is that the structure is in a sense ‘minimal’ and therefore easier to work with in concrete examples. [^3] The model-complete core of ${\mathfrak{B}}_{\mu}$ would be a natural choice for the canonical dual of $\mu$ to work with instead of ${\mathfrak{B}}_{\mu}$. However, proving that the model-complete core has a finitely bounded homogeneous expansion (so that, for example, Theorem 3.4 applies) requires introducing further model-theoretical notions (Mottet and Pinsker, 2021) which we want to avoid in this article.
Figure 5. Visualization of the query $\mu$ from (18).:
@@
There is a finitely bounded homogeneous dual $\mathfrak{B}$} of $\mu$ such that the valued $\tau$-structure $\Gamma:=\Gamma({\mathfrak{B}},\emptyset)$ has a binary fractional polymorphism which is canonical and pseudo cyclic with respect to $\operatorname{Aut}({\mathfrak{B}})=\operatorname{Aut}(\Gamma)$. Hence, $\operatorname{VCSP}(\Gamma)$ and the resilience problem for $\mu$ are in P. The polynomial-time tractability result even holds for resilience of $\mu$ with exogenous relations from any $\sigma\subseteq\tau$.
The rest of this section is devoted to the proof of Proposition 8.29.
Figure 6. An illustration of a finite substructure of $\mathfrak{B}$} from the proof of Proposition 8.29 that contains representatives for all orbits of pairs of $\operatorname{Aut}({\mathfrak{B}})$. Arrows are not drawn on undirected edges.:
Since the Gaifman graph of $\mu$ is a complete graph, there exists the structure ${\mathfrak{H}}_{\mu}$ as in Theorem 8.19. Let $\mathfrak{B}$} be the model-complete core of ${\mathfrak{H}}_{\mu}$. Note that $\mathfrak{B}$} has the property that a countable structure $\mathfrak{A}$} maps homomorphically to $\mathfrak{B}$} if and only if ${\mathfrak{A}}\models\neg\mu$; in particular, $\mathfrak{B}$} is a dual of $\mu$ and ${\mathfrak{B}}\models\neg\mu$. The structure ${\mathfrak{H}}_{\mu}$ is homogeneous, and it is known that the model-complete core of a homogeneous structure is again homogeneous (see Proposition 4.7.7 in (Bodirsky, 2021)), so $\mathfrak{B}$} is homogeneous. Let $\Gamma:=\Gamma({\mathfrak{B}},\emptyset)$.
@@
Suppose for contradiction that (21) does not hold and let $x$ and $y$ be witnesses, in particular, $x\neq y$. Then $\neg S(x)\vee\neg R(y,y)$ and $\neg R(x,x)\vee\neg S(y)$, i.e., $\neg S(x)\wedge\neg R(x,x)$, or $\neg S(x)\wedge\neg S(y)$, or $\neg R(y,y)\wedge\neg R(x,x)$, or $\neg R(y,y)\wedge\neg S(y)$. In each of these cases we may add both $R$-edges between the distinct elements $x$ and $y$ to $\mathfrak{B}$} and obtain a structure not satisfying $\mu$, which leads to a contradiction as above.
For an illustration of a finite substructure of $\mathfrak{B}$} which contains a representative for every orbit of pairs in $\operatorname{Aut}({\mathfrak{B}})$; see Figure 6. A homomorphism is called strong if it also preserves the complements of all relations (note that an injective strong homomorphism is an embedding).
Claim 8.30.
For every finite $\tau$-structure $\mathfrak{A}$} that satisfies $\neg\mu$ and the sentences in (20) and (21), there exists a strong homomorphism to $\mathfrak{B}$}.
Proof.
We first observe that $\mathfrak{B}$} embeds the relational $\tau$-structure $\mathfrak{A}$$^{\prime}$ on the domain $A^{\prime}=\mathbb{N}\cup\{a\}\cup\{b\}$, where $a,b\notin\mathbb{N}$ are distinct, $S^{{\mathfrak{A}}^{\prime}}=\{a\}$ and
@@
To see this, first note that there exists a homomorphism $h\colon{\mathfrak{A}}^{\prime}\to{\mathfrak{B}}$, because ${\mathfrak{A}}^{\prime}\models\neg\mu$. If $h$ is not an embedding, then $h$ is not a strong homomorphism or it is not injective. Note that if ${\mathfrak{B}}\models S(h(x))$ for $x\in\mathbb{N}\cup\{b\}$, then ${\mathfrak{B}}\models\mu$, because $h$ is a homomorphism. Similarly we get ${\mathfrak{B}}\models\mu$ if ${\mathfrak{B}}\models R(h(a),h(b))$ or ${\mathfrak{B}}\models R(h(x),h(x))$ for $x\in\mathbb{N}\cup\{a\}$. Finally, if $h(x)=h(y)$ for $x,y\in A^{\prime}$, we can again verify that the substructure induced by $h(A^{\prime})$ in $\mathfrak{B}$} satisfies $\mu$. In each of these cases, we get a contradiction with $\mathfrak{B}$} being a dual of $\mu$. Therefore, $\mathfrak{A}$$^{\prime}$ embeds into $\mathfrak{B}$}. In particular, there are infinitely many $x\in B$ such that ${\mathfrak{B}}\models\neg S(x)\wedge\neg R(x,x)$ and by (21), for every $y\in B$, $x\neq y$, we have ${\mathfrak{B}}\models R(x,y)\wedge R(y,x)$.
@@
Let $h$ be a homomorphism from $\mathfrak{A}$} to $\mathfrak{B}$}, which exists since ${\mathfrak{A}}\models\neg\mu$. If $s(h)+r(h)=0$, then $h$ is a strong homomorphism and there is nothing to prove. Suppose therefore $s(h)+r(h)>0$, equivalently, $s(h)>0$ or $r(h)>0$. We construct a homomorphism $h^{\prime}$ such that $r(h^{\prime})+s(h^{\prime})<r(h)+s(h)$. Since $r(h)+s(h)$ is finite, by applying this construction finitely many times, we obtain a strong homomorphism from $\mathfrak{A}$} to $\mathfrak{B}$}.
First suppose that $s(h)>0$. Then there exists $a\in A\setminus S^{{\mathfrak{A}}}$ such that $h(a)\in S^{{\mathfrak{B}}}$. By (19), ${\mathfrak{B}}\not\models R(h(a),h(a))$ and hence ${\mathfrak{A}}\not\models R(a,a)$. Pick $b\in B\setminus h(A)$ such that ${\mathfrak{B}}\models\neg S(b)\wedge\neg R(b,b)$ (recall that there are infinitely many such $b\in B$) and define
@@
Observe that $h^{\prime}$ is a homomorphism, $s(h^{\prime})<s(h)$ and $r(h^{\prime})=r(h)$.
Second suppose that $r(h)>0$. Then there exists $(x,y)\in A^{2}\setminus R^{{\mathfrak{A}}}$ such that $(h(x),h(y))\in R^{{\mathfrak{B}}}$. If $x=y$, the argument is similar as in the case $s(h)>0$. Finally, if $x\neq y$, then ${\mathfrak{A}}\models(S(x)\wedge R(y,y))\vee(R(x,x)\wedge S(y))$, because $\mathfrak{A}$} satisfies the sentence in (21). Since $\mathfrak{A}$} satisfies the sentence in (20), ${\mathfrak{A}}\models R(y,x)$. Since $h$ is a homomorphism, we have
@@
which contradicts ${\mathfrak{B}}\not\models\mu$. ∎
Claim 8.31.
Every finite $\tau$-structure $\mathfrak{A}$} that satisfies $\neg\mu$ and the universal sentences in (20) and (21) embeds into $\mathfrak{B}$}. In particular, $\mathfrak{B}$} is finitely bounded.
Proof.
Let $\mathfrak{A}$} be such a structure. By Theorem 8.19, there is an embedding $e$ of $\mathfrak{A}$} into ${\mathfrak{H}}_{\mu}$. Since ${\mathfrak{H}}_{\mu}$ is homogeneous and embeds every finite $\tau$-structure that satisfies $\neg\mu$, there exists a finite substructure $\mathfrak{A}$$^{\prime}$ of ${\mathfrak{H}}_{\mu}$ satisfying the sentences in (20) and (21) such that $e({\mathfrak{A}})$ is a substructure of $\mathfrak{A}$$^{\prime}$ and for all distinct $a,b\in A$ there exists $s\in S^{{\mathfrak{A}}^{\prime}}$ such that ${\mathfrak{H}}_{\mu}\models R(e(a),s)\wedge R(s,e(b))$. By Claim 8.30, there is a strong homomorphism $h$ from $\mathfrak{A}$$^{\prime}$ to $\mathfrak{B}$}.
We claim that $h\circ e$ is injective and therefore an embedding of $\mathfrak{A}$} into $\mathfrak{B}$}. Suppose there exist distinct $a,b\in A$ such that $h(e(a))=h(e(b))$. Since $e({\mathfrak{A}})$ satisfies the sentence in (20), we have ${\mathfrak{B}}\models R(h(e(a)),h(e(a)))$. Let $s\in S^{{\mathfrak{A}}^{\prime}}$ be such that ${\mathfrak{H}}_{\mu}\models R(e(a),s)\wedge R(s,e(b))$. Hence,
@@
a contradiction to ${\mathfrak{B}}\not\models\mu$. It follows that $h\circ e$ is an embedding of $\mathfrak{A}$} into $\mathfrak{B}$}. ∎
@@
(22) ${\mathfrak{M}},{\mathfrak{N}}$ $\models R\big((x_{1},y_{1}),(x_{2},y_{2})\big)$ $\text{ if }{\mathfrak{B}}$ $\models R(x_{1},x_{2})\wedge R(y_{1},y_{2}),$
(23) ${\mathfrak{M}},{\mathfrak{N}}$ $\models S\big((x,y)\big)$ $\text{ if }{\mathfrak{B}}$ $\models S(x)\wedge S(y)$
(24) ${\mathfrak{M}}$ $\models S\big((x,y)\big)$ $\text{ if }{\mathfrak{B}}$ $\models S(x)\vee S(y)$
(25) ${\mathfrak{N}}$ $\models R\big((x,y),(x,y)\big)$ $\text{ if }{\mathfrak{B}}$ $\models R(x,x)\vee R(y,y).$
@@
Figure 7. An illustration of the conditions (A) and (B) in $\mathfrak{M}$} and $\mathfrak{N}$}.:
Claim 8.32.
$\mathfrak{M}$} and $\mathfrak{N}$} satisfy the sentence in (20).
Proof.
We prove the statement for $\mathfrak{M}$}; the proof for $\mathfrak{N}$} is similar. Let $(x_{1},y_{1}),(x_{2},y_{2})\in B$ be such that $(x_{1},y_{1})\neq(x_{2},y_{2})$ and ${\mathfrak{M}}\models\neg R((x_{2},y_{2}),(x_{1},y_{1}))$. Since $\mathfrak{M}$} satisfies the sentence in (21), we must have either that ${\mathfrak{M}}\models S(x_{1},y_{1})\wedge R((x_{2},y_{2}),(x_{2},y_{2}))$ or that ${\mathfrak{M}}\models S(x_{2},y_{2})\wedge R((x_{1},y_{1}),(x_{1},y_{1}))$. Suppose the former is true; the other case is treated analogously. Then ${\mathfrak{B}}\models R(x_{2},x_{2})\wedge R(y_{2},y_{2})$ and ${\mathfrak{B}}\models S(x_{1})\vee S(y_{1})$. If ${\mathfrak{B}}\models S(x_{1})$, then $x_{1}\neq x_{2}$ and by (20) we have ${\mathfrak{B}}\models R(x_{1},x_{2})\vee R(x_{2},x_{1})$. By (20) and (21) for $(y_{1},y_{2})$, we obtain that ${\mathfrak{M}}\models R((x_{1},y_{1}),(x_{2},y_{2}))$ by (22) or one of the conditions (A)-(D). The argument if ${\mathfrak{B}}\models S(y_{1})$ is similar with $x$ and $y$ switched. ∎
Claim 8.33.
$\mathfrak{M}$} and $\mathfrak{N}$} satisfy $\neg\mu$. Let $x_{1},x_{2},y_{1},y_{2}\in B$.
Proof.
Suppose for contradiction that